OpenStudy (anonymous):

The sum of -4x and 2 is 9 ?

OpenStudy (anonymous):

-4x+2=9 SO we need to solve?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-4x+2=9 -4x=7 x=-7/4

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yep thanks you very much :)